OpenStudy (anonymous):

Simplify: ( y^−7 × y^−3 )^−1

OpenStudy (anonymous):

we should sum powers of variables so write in this form: (y^-7)(y^-3)=y^(-7-3)=y^-10 and then ..step 2 multiply powers: (y^-10)^(-1)=y^((-10)(-1))=y^10